Suppose that Carl was riding a bicycle at 15 mph, he rode 10 miles farther than Michael, who was riding his bike at 14 mph. Carl rode for 30 minutes longer than Michael. How long did Carl ride his
bike?

Answer by stanbon(57301)

Suppose that Carl was riding a bicycle at 15 mph, he rode 10 miles farther than Michael, who was riding his bike at 14 mph. Carl rode for 30 minutes longer than Michael. How long did Carl ride his bike?
-------------------
Carl DATA:
rate = 15 mph ; time = x hrs; distance = 15x miles
-------------------
Michael DATA:
rate = 14 mph ; time = x-(1/2) hrs: distance = 14(x-(1/2)= 14x-7 miles
-------------------------
Equation:
Carl distance - Michael distance = 10 miles
--
15x - (14x-7) = 10
15x-14x+7 = 10
x = 3 hrs (time Carl rode his bike)
======================================
Cheers,
Stan H.